Grade 12Algebra

Give first the step you will use to separate the variable and then solve the equation:
x + 1 = 0

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to subtract 1 to both the side of given equation,
Then we get,
= x +1 - 1 = 0 - 1
= x = - 1

Thanks